# Christian Peukert [^1] Corresponding author. Florian Abeillon Jérémie Haese Franziska Kaiser Alexander Staub

### v1

@@ 5.3.1 Nature-themed and curated images
Note: The dependent variable is an indicator of whether a user account has uploaded at least one image in a given month. Treated indicates whether a user was part of the LITE dataset. Fixed effects for upload month of the image and user account. Standard errors clustered at the user level in parentheses. $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$
@@
Next, we study whether the types of images uploaded to Unsplash have changed. In column (1) of Table 6, we show that the likelihood of uploading at least one nature-themed image decreases by 23 percentage points or 30% relative to the average in the before period. Note that this is about the same effect size we get when looking at all types of images in column (2) of Table 3. However, columns (2) and (3) of Table 6, suggest that the extensive margin decreases at 60% for curated images and at 70% for nature-themed and curated images. If we believe that Unsplash’s curation team picks high-quality images, then this suggests a reduction in the artistic quality of uploads.
5.3.2 Analysis of Variety and Novelty measures
Figure 2: Changes in uploads, similarity of images
Panel A: Levels of average similarity
Panel B: Differences in average similarity
Panel C: Levels of very similar count
Panel D: Differences in very similar count
Note: In Panel A, we depict the average similarity of uploaded images by users in the treatment group (solid) and by users in the control group (dashed). In Panel C, we plot the number of images uploaded by treated users (solid) and control users (dashed) that are very similar to the existing stock (threshold 0.8). In Panels B and D, we plot OLS estimates of the $\delta_{\tau}$ coefficients are obtained from $Y_{it}=\sum_{\tau\in T}\nu_{\tau}+\sum_{\tau\in T}\delta_{\tau}\left(\gamma_{\tau}\times Treated_{i}\right)+\varepsilon_{it}$, with the number of uploads of user $i$ per month $t$ as the dependent variable. The dots reflect month-specific point estimates comparing treatment users to control users. Standard errors are clustered at the user level, and bars indicate 90% confidence bands. In all panels, the vertical line indicates the creation of LITE 1.0.0, and numbers on the horizontal axis indicate the number of months before and after that.
Table 7: Results: Changes in uploads, similarity of images
Avg. Similarity LogVerySimilar (0.8) LogVerySimilar (0.9)
(1) (2) (3) (4) (5) (6)
(0.00097) (0.00345) (0.01817) (0.05707) (0.02300) (0.06871)
Mean Treated Before 0.3910 0.3910 8.1285 8.1285 4.5527 4.5527
Image-Age FE Yes Yes Yes Yes Yes Yes
User FE No Yes No Yes No Yes
Observations 340,991 340,346 340,991 340,346 340,991 340,346
Note: The dependent variable in columns (1) and (2) measures how similar an uploaded image is relative to the average image in the stock of images uploaded by treated vs. control users up until one year before the creation of the LITE dataset. In columns (3)–(6), the dependent variable is the log+1 count of images from the stock of images uploaded by treated/control users up until one year before the creation of the LITE dataset that have a cosine similarity score of larger than 0.8 or 0.9 relative an uploaded image. Treated indicates whether a user was part of the LITE dataset. Fixed effects for upload month of the image and user account in columns (2), (4) and (6). Standard errors in parentheses, White-robust in columns (1), (3) and (5), and clustered at the user level in columns (2), (4) and (6). $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$
Next, we study how strategic behavior affects the variety and novelty of contributions more directly. In Table 7, we report results from our analysis of similarity using natural language processing of user-specified keywords of uploaded images. We use the same difference-in-differences identification strategy as above. Panel A of Figure 2 shows that the unconditional mean of the average similarity to the pre-existing stock of photos decreases less for images uploaded by treated users than for images uploaded by control users. In other words, treated users upload images that are less novel and more similar to the existing stock of images. This is confirmed more formally in Panel B, where we plot estimates of the difference between the treatment and control group in each month, before and after the release of the LITE dataset, from a model without user-fixed effects. In Panels C and D, we turn to a count of the number of very similar images. Again, we see that the decrease in very similar images uploaded by users in the treatment group is less steep than in the control group.
We quantify these differences in Table 7. We find that it is not users who change the type of images they upload, but a difference in the type of uploaded image is coming from a change in the composition of users. In the specifications with user-fixed effects in columns (2), (4) and (6), we do not find much evidence that the release of the LITE dataset caused new uploads to differ from existing uploads. This is especially the case concerning the number of highly similar images in the existing stock. Hence, users do not change the types of images they upload. However, looking at the results from specifications without user-fixed effects in columns (1), (3) and (5), we do see that uploaded images become more similar to the existing stock. Column (1) suggests an increase of about 5% in average similarity for uploads by treated users, whereas columns (3) and (5) suggest an increase of about 28% and 13%, respectively, in the number of highly similar images. Overall, the results imply that the composition of the training dataset will change because of changes in the user composition and the types of images they upload, not because individual users change the types of images they upload. This is in line with the results we describe above. In what follows, we explore potential mechanisms that may explain the changes in user behavior.
5.4 Mechanisms: Awareness of AI potential and treatment intensity
Table 8: Results: Changes in uploads, mechanisms
Uploads
(1) (2)
Post $\times$ Treated -1.0056∗∗∗
PostAug22 $\times$ Treated -0.4054∗∗∗
Post $\times$ TreatedSingle -0.3059∗∗∗
Post $\times$ TreatedMultiple -2.2830∗∗∗
Mean Treated Before 2.9038
Mean TreatedSingle Before 1.4011
Mean TreatedMultiple Before 5.0982
Month FE Yes Yes
User FE Yes Yes
Observations 612,662 612,662
Note: The dependent variable is the total number of uploads per month. Post indicates the entire period after the creation of LITE and PostAug22 indicates the period after August 2022, when groundbreaking genAI for images was released. Treated indicates whether a user was part of the LITE dataset. TreatedSingle indicates whether a user had exactly one image in the LITE dataset. TreatedMultiple indicates whether a user had more than one image in the LITE dataset. Fixed effects for upload month of the image and user account. Standard errors clustered at the user level in parentheses. $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$
The results in Table 8 serve to highlight potential mechanisms through which the inclusion in the LITE dataset induces users to reduce their upload activity. In column (1), informed by Figure 1, we split the post-period in two parts. The first 24 months correspond to the period between June 2020 and August 2022, a period with modest public interest in AI and AI art in particular. After August 2022, however, there was a substantial increase in media coverage on the topic, e.g., a New York Times article titled “We Need to Talk About How Good A.I. Is Getting”, which may have further raised awareness about genAI in the Unsplash community.202020see https://www.nytimes.com/2022/08/24/technology/ai-technology-progress.html The timing also coincides with the public release of Stable Diffusion, the first text-to-image generator that could be run on consumer-grade machines, democratizing access to genAI technology after similar technology released a few months earlier had only restricted public access (e.g. DALL-E and Midjourney). [^21] See https://stability.ai/news/stable-diffusion-announcement. We find a treatment effect of 34% before August 2022 and 49% thereafter, which corresponds to a significant 40% increase in the treatment effect. This suggests that, as awareness of the commercial potential of image generation AI raises, photographers reduce their contributions to Unsplash and, therefore, to the flow of AI training data.
In column (2), we distinguish between users that were affected more or less intensively by the release of the LITE dataset. In particular, we distinguish between users who had one image included (about 60%) and users who had multiple images included (about 40%). Conditional on more than one image in the LITE dataset, the average number of images per user is 5.96 (standard deviation 14.18 and maximum 419). The estimated effect size is 22% for single-treated users and 45% for multiple-treated users. This difference is statistically significant. These results provide further evidence that users must have been aware of their images being included in the LITE dataset because otherwise, we would not expect a stronger response from more heavily treated users. This result has interesting implications for policy, suggesting that a one-size-fits-all approach to licensing of human-generated work for AI training datasets can lead to unintended heterogeneous effects on the flow of training data. As a consequence, our results suggest that a licensing scheme that ensures a continued flow of data may need to compensate more active contributors to a larger degree (Wang et al., 2024).
Finally, it is worth noting that we observe that users in the treatment group are three times more likely to be contributors in the Unsplash+ program (t-statistic in a two-sided t-test is 6.13). This program enables users to gain monetary compensation and explicitly disallows the use of their works for AI training. Hence, we interpret the substantially higher uptake of Unsplash+ among users who had at least one image in the LITE dataset as a strategic response to counteract economic forces and to avoid the utilization of their works in AI applications.
5.5 Robustness
5.5.1 Alternative control group definition
Table A.1 repeats the specification of our baseline estimations in Table 3 but with different definitions of the control group. In columns (1)–(3), we report results using a control group composed of users that did not have nature-themed (i.e., not with the keyword “nature”) and curated images at the time when the LITE dataset was created. The effect size for the number of uploads is 47% in column (1), 33% in terms of the extensive margin in column (2), and 19% when using a specification with the log-transformed number of uploads in column (3). In columns (4)–(6), we use a control group comprised of all users whose images were not included in the LITE dataset. The effect sizes are 50%, 40% and 21%, respectively. Note that the estimates reported in Table A.1 are not significantly different from those reported in Table 3 in the sense that 90% confidence bands overlap.
5.5.2 OLS versus non-linear models
Given that our outcome variable is both a count, i.e. a non-negative integer, and zero in 86.6% of the observations, one might be worried that OLS is not the optimal choice of estimator (Wooldridge, 2023). Table A.2 shows that our baseline results are broadly robust to using a Poisson pseudo-maximum likelihood regression (PPML) approach. [^22] We do not report results from a negative binomial model because PPML is preferred for its robustness (Blackburn, 2015). However, the results indicate that PPML is sensitive to large values of the DV, which is of economic interest in our setting (i.e. heavy platform users). In addition, the interaction effect in non-linear models cannot be directly interpreted as the average treatment effect on the treated (Ai and Norton, 2003; Puhani, 2012) while simultaneously imposing additional restrictions for parallel trends to hold (Roth and Sant’Anna, 2023). For these reasons, and for simplicity in the interpretation of effect sizes, we report estimates from OLS regressions throughout.
6 Discussion
6.1 Back-of-the-envelope estimation of aggregate effects
Having estimated the average treatment effects following the release of the LITE dataset on several key outcomes, we can now use these estimations to simulate several counterfactuals in which we vary the proportion of users treated.
To proceed, we first need to define a few key variables. The share of users included in the LITE dataset is $\lambda$, the effect of being in the LITE dataset is $\hat{\delta}$, and the factual growth rate of the number of images per user in the post-period is $g$. With this, we can calculate a counterfactual growth rate $\bar{g}=g/(1-\lambda+\lambda\hat{\delta})$, which is the growth rate that would have materialized had the LITE dataset not been released. Then, we can use a few descriptive statistics as well as our causal estimates to attempt a back-of-the-envelope calculation of counterfactuals of key outcomes such as the total number of uploads. Finally, once we have calculated $\bar{g}$, we can simulate any growth rate $\tilde{g}(\lambda)=(1-\lambda)(1+\bar{g})+\lambda(1+\hat{\delta}\bar{g})-1$. This is useful to calculate counterfactuals where we can vary the size of the LITE training dataset and approximate how the size of the LITE training dataset affects the flow of data in terms of the aggregate number of data points added to Unsplash.
To simulate the impact of the share of treated users on the number of uploads, we use the estimated average treatment effect on the treated and the mean uploads of treated users from Table 3. This leads to $\hat{\delta}=1-(1.1099/2.9038)=0.6$. When the LITE dataset was created in June 2020, Unsplash had a stock of $S_{0}=1,931,324$ images uploaded by $U_{0}=198,792$ users. The LITE dataset encompasses $U_{L}=8,298$ users, such that the share of users in the LITE dataset was $\lambda=U_{L}/U_{0}=0.042$. About three years later, in May 2023, Unsplash has $S_{t}=4,960,120$ images and $U_{t}=338,635$ users. Hence, the factual growth rate $g=\frac{S_{t}}{U_{t}}/\frac{S_{0}}{U_{0}}-1=14.65/9.72-1=0.51$ and therefore the counterfactual growth rate $\bar{g}=(1.51)/(1-0.042+0.042\times 0.6)=0.52$. In absolute terms, without the release of the LITE dataset, users would have uploaded 19,313 more images to Unsplash ($=S_{0}(0.52-0.51)$).
We can now extrapolate that a twice-as-large LITE dataset would have reduced the flow of data by 2% (=1-0.50/0.51) compared to the factual and making the entire catalog available for commercial AI research would have reduced the flow of data by 39% (=1-0.31/0.51), corresponding to 386,264 images uploaded less over the period.
6.2 How different is Unsplash from other platforms?
To provide some insight into this question, it is useful to compare Unsplash to Instagram, the by far largest image-sharing platform. According to Instagram’s terms of service, users retain ownership of any intellectual property rights that they hold in that content. However, by posting content to Instagram, a user grants the platform a non-exclusive, royalty-free, transferable, sublicensable, worldwide license to use the content they post. [^23] See https://help.instagram.com/478745558852511/. This means Instagram can use, reproduce, modify, display, and distribute the content, and they can also allow others to do so within their platform or in association with other media. This license ends when a user deletes the content or the account unless the content has been shared with others and they have not deleted it. Unsplash operates in a similar fashion yet in an even more permissive fashion. When a user uploads an image to Unsplash, they grant anyone a license to use their image for free, including for commercial purposes, without requiring permission from or providing attribution to the user. [^24] See https://unsplash.com/license. By uploading photos to Unsplash, users agree to this license, which effectively allows anyone worldwide to use the images with very few restrictions. It is important to note that while a user still retains the copyright to their image, the rights they grant through the Unsplash license are much broader in scope and do not end even if they remove the image from the platform. A key difference there is the control over use. On Instagram, the use of an uploaded image is mostly within the context of the platform, though the license a user grants Instagram is quite broad. Unsplash, in contrast, allows anyone to use the images for nearly any purpose, including commercially and outside the platform, without any compensation or credit to the uploader. Instagram’s terms are not explicitly clear about commercial use by third parties unless it is sublicensed, while Unsplash explicitly allows commercial use by anyone.
@@ 6.2 How different is Unsplash from other platforms?
Note: The dependent variable is the number of monthly uploads to either Unsplash or Instagram, or both. Treated indicates whether a user was part of the LITE dataset. Standard errors clustered at the user level in parentheses. $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$
@@
In Table 9, we study whether users’ reaction to having their images included in an AI training dataset affects only their uploads to Unsplash, or whether they also change their behavior on Instagram. To do so, we compare the number of uploads of treated versus control users on Unsplash to their uploads on Instagram. In column (1), we see that the average treatment effect with respect to uploads on Unsplash is about half the size of our baseline (column 1 in Table 3) in this restricted sample. In column (2), we do not find a significant effect on Instagram uploads, as the point estimate is very close to zero. In columns (3) and (4), we combine information from both platforms to compare effect sizes directly. The linear combination in both sets of results, with user fixed effects and with user-platform fixed effects, again indicates that the effect on Instagram is not significant. This implies that users did not switch to Instagram as a new outlet for their photography work, but it also does not imply that users stopped to publish their work anywhere online. Overall, our results suggest that adding Unsplash images to an Unsplash-curated AI training dataset affects user behavior on Unsplash, but not elsewhere.

### v2

@@ 4.2.4 Differentiation or exit?
Note: The dependent variable is an indicator of whether a user account has uploaded at least one image in a given month. Column (1) includes only Nature-themed images, column (2) includes only Curated images, and column (3) focuses on images that are both Nature-themed and Curated.Treated indicates whether a user was part of the LITE dataset. Fixed effects for upload month of the image and user account. Standard errors are clustered at the user level in parentheses. $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$
@@
Next, we study whether contributors differentiate and upload different types of images to Unsplash. In column (1) of Table 6, we show that the likelihood of uploading at least one nature-themed image decreases by 23 percentage points or 30% relative to the average in the before period. Note that this resembles approximately the same effect size we get when looking at all types of images in column (2) of Table 3. However, columns (2) and (3) of Table 6 suggest that the extensive margin decreases at 60% for curated images and at 70% for nature-themed and curated images. If we believe that Unsplash’s curation team picks high-quality works, then this suggests a reduction in the artistic quality of uploads, especially in the nature genre.
@@
Note: In column (1), the dependent variable is the number of monthly uploads to Unsplash. In column (2), the dependent variable is the number of monthly uploads to Instagram, and in column (3), the dependent variable is the number of uploads to both platforms. Treated indicates whether a user was part of the LITE dataset. Standard errors are clustered at the user level in parentheses. $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$
The next question is whether the behavioral changes of treated users have broader consequences or are limited to Unsplash. In other words, do photographers who leave Unsplash, reduce their uploads or lower image quality also alter their professional approach or exit the profession entirely, or are their reactions solely a response to perceived unfair treatment by the platform?
@@
To answer this question, we study whether treated users only reduce their contributions to Unsplash or another image-sharing platform as well. As described in section 4.1.4, for a subset of users, we have access to their upload activities on both Unsplash and Instagram, before and after Unsplash releases data. Panel A (showing upload levels) and Panel C (showing the difference in uploads between Instagram and Unsplash within the treatment group) of Figure 2 suggest that inclusion in the LITE dataset affects only activity on Unsplash, with no spillovers to Instagram. In Table 7, we provide the corresponding econometric results. First, we see in column (1) that the average treatment effect concerning uploads on Unsplash is about half the size of our baseline (column 1 in Table 3) in the subsample of users active on Unsplash and Instagram. In column (2), we do not find a significant effect on Instagram uploads and the point estimate is very close to zero. In columns (3) and (4), we combine information from both platforms to compare effect sizes directly. The linear combination in both sets of results, with user fixed effects and with user-platform fixed effects, again indicates that the impact on Instagram is not significant. This implies that users did not switch to Instagram as a new outlet for their photography work, but it does not imply that users stopped publishing their work anywhere online. Overall, our results suggest that adding Unsplash images to an Unsplash-curated AI training dataset affects user behavior on Unsplash, but not elsewhere. More broadly, this leads to the conclusion that the program likely did not incentivize photographers to exit their profession.
4.3 Implications for AI training data
4.3.1 Analysis of variety and novelty measures
Figure 3: Examples of images based on variety and novelty scores
High variety
High novelty
High variety
High novelty
Mean variety
Mean novelty
Low variety
Low novelty
Low variety
Low novelty
Note: These examples illustrate how images vary in both variety and novelty according to our four-step scoring procedure. Variety reflects the average similarity across all existing images (high average similarity means low variety score), whereas novelty captures the count of images whose similarity to the new upload exceeds a specified threshold (high count means low novelty score). The first two images (left) exhibit relatively high variety and novelty, the middle image represents mean variety and novelty, and the final two images (right) show low variety and novelty. All examples are drawn from the Unsplash sample as of May 2023. Image credits (from left to right): Ashe Walker, Wolfgang Hasselmann, Emre, Marek Piwnicki, and Engin Akyurt.
We now turn our attention to the image level. To investigate whether contributions to Unsplash change after the release of the LITE dataset, we compare the flow of new uploads to the stock of images one year before the launch of the research program. In particular, to capture variety as a quality measure of a training dataset, we measure how similar each new upload is relative to all existing images. To capture novelty, we count the number of images in the stock of data that are very similar to a focal new upload.
We then compare how variety and novelty evolve over time for treated users and control users, before and after the release of the LITE dataset. We do so based on the keywords that users assign to their uploaded images, which on average gives us 10.28 keywords per image. To assess the similarity between images we proceed in four steps of NLP.
In step one, we aggregate keywords for each image and train a Word2Vec model on these aggregated keywords to generate a vector representation for each unique keyword. Herein, the model is configured to create 100-dimensional vectors for each keyword, with a context window of 5 words and a minimum word frequency threshold of 1, ensuring that even rare keywords are included in the model. In step two, we apply the Word2Vec model to each image, converting the keywords into vectors, and then calculate the average vector to represent the image’s semantic content. In step three, we calculate cosine similarity scores between the vector representation of a particular image and the vector representations of all existing images. Finally, in step four, we compute the average similarity across all images and count the number of images that exceed a specified similarity threshold, respectively.
We perform steps 1 and 2 for the stock of images uploaded by users that will be treated eventually as well as for the stock of images uploaded by users in the control group. We define the stock as all images uploaded before June 25, 2019. We perform steps 1-4 for all images uploaded by treated and control users after June 25, 2019. This allows us to compare new uploads starting one year before the start of Unsplash’s research program to new uploads after the release of the first LITE dataset, across users in the treatment and in the control group. In step 4, variety is defined as the average cosine similarity between a new upload and the images in the stock, while novelty is defined as the count of stock images whose similarity to the new upload exceeds a threshold (here 0.8 or 0.9). Formally, if $s(\mathbf{x}_{i},\mathbf{x}_{j})$ is the cosine similarity between the embeddings of images $i$ and $j$, and $\theta$ is a chosen threshold, then the novelty measure for image $i$ is:
$\sum_{j\in\text{Stock}}\mathbb{I}\bigl{[}s(\mathbf{x}_{i},\mathbf{x}_{j})>\theta\bigr{]},$
where $\mathbb{I}[\cdot]$ is an indicator function. A higher count indicates that $i$ has more “near neighbors” in the stock and is thus less novel. Conversely, a smaller count implies greater novelty. Figure 3 presents five images illustrating different segments of the variety–novelty space. From left to right, the first two show images with high variety and high novelty scores, the middle image represents mean variety and novelty scores, and the final two exhibit low variety and low novelty scores. This highlights how images can vary in their average similarity to the stock (variety) and the number of “near neighbors” above a similarity threshold (novelty).
Figure 4: Changes in uploads, similarity of images
A: Variety - Levels of average similarity
B: Variety - Differences in average similarity
C: Novelty - Levels of number of very similar images
D: Novelty - Differences in number of very similar images
Note: In Panel A, we depict the demeaned average similarity of uploaded images by users in the treatment group (solid) and by users in the control group (dashed). In Panel C, we plot the demeaned number of images uploaded by treated users (solid) and control users (dashed) that are very similar to the existing stock (threshold 0.8). We use the group-specific pre-period mean number of uploaded images to demean (0.38 and 8,403 for the treatment group, and 0.39 and 8,566 for the control group). In Panels B and D, we plot OLS estimates of the $\delta_{\tau}$ coefficients obtained from estimating $Y_{it}=\sum_{\tau\in T}\nu_{\tau}+\sum_{\tau\in T}\delta_{\tau}\left(\gamma_{\tau}\times Treated_{i}\right)+\varepsilon_{it}$, with the number of uploads of user $i$ per month $t$ as the dependent variable. The dots reflect month-specific point estimates comparing treatment users to control users. Standard errors are clustered at the user level, and bars indicate 90% confidence bands. In all panels, the vertical line indicates the creation of LITE 1.0.0, and numbers on the horizontal axis indicate the number of months before and after that.
Table 8: Results: Changes in uploads, similarity of images
Avg. Similarity LogVerySimilar (0.8) LogVerySimilar (0.9)
(1) (2) (3) (4) (5) (6)
(0.00097) (0.00345) (0.01817) (0.05707) (0.02300) (0.06871)
Mean Treated Before 0.3910 0.3910 8.1285 8.1285 4.5527 4.5527
Image-Age FE Yes Yes Yes Yes Yes Yes
User FE No Yes No Yes No Yes
Observations 340,991 340,346 340,991 340,346 340,991 340,346
Note: The dependent variable in columns (1) and (2) is computed as the average similarity between an uploaded image and the stock of images uploaded by treated vs. control users up until one year before the creation of the LITE dataset. In columns (3)–(6), the dependent variable is the log+1 count of images from the stock of images uploaded by treated/control users up until one year before the creation of the LITE dataset that have a cosine similarity score of larger than 0.8 or 0.9 relative to an uploaded image. Treated indicates whether a user was part of the LITE dataset. Fixed effects for upload month of the image and user account in columns (2), (4), and (6). Standard errors in parentheses, White-robust in columns (1), (3), and (5), and clustered at the user level in columns (2), (4), and (6). $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$
Panel A of Figure 4 shows that the demeaned average similarity to the pre-existing stock of images decreases less for images uploaded by treated users than for images uploaded by control users. In other words, treated users upload works that vary less and are more similar to the existing stock of images. This is confirmed more formally in Panel B, where we plot estimates of the difference between the treatment and control group in each month, before and after the release of the LITE dataset, from a model without user-fixed effects. In Panels C and D, we turn to our novelty measure, a count of the number of very similar images. Again, we see that the decrease in very similar images uploaded by users in the treatment group is less steep than in the control group. Finally, the lack of statistically significant deviations from 0 in the pre-period (all month-specific differences are jointly not significantly different from zero) supports the parallel trends assumption of the difference-in-differences model.
We quantify these differences in Table 8. We find that it is not users who change the type of images they upload, but the difference in the type of uploaded image derives from a change in the composition of users. In the specifications with user-fixed effects in columns (2), (4) and (6), we do not find much evidence that the release of the LITE dataset caused new uploads to differ from existing uploads. This is especially the case concerning the number of highly similar images in the existing stock. Hence, users do not change the types of images they upload. However, looking at the results from specifications without user-fixed effects in columns (1), (3), and (5), we do see that uploaded images become more similar to the existing stock. Column (1) suggests an increase of about 5% in average similarity for uploads by treated users, whereas columns (3) and (5) suggest an increase of about 29% and 14%, respectively, in the number of highly similar images.
4.3.2 Link to AI output quality
Having quantified how user behavior changes the quantity, variety, and novelty of the flow of data, we now discuss the potential knock-on effects on AI output quality. First, ample evidence suggests that algorithmic performance is positively related to dataset size (see e.g., Budach et al., 2022; Valavi et al., 2022b; Peukert et al., 2024). This corresponds to evidence showing that the AI industry is vying to secure access to big datasets (Beraja et al., 2023). Although returns to scale are decreasing, a rough linear approximation of the estimates suggests that a 2% reduction in sample size, like the one we find in our context, implies a 1.4–2% reduction in algorithmic performance (see Figure 2 in Peukert et al., 2024 and Figure 12 in Budach et al., 2022).
Second, regarding the quality of data, a growing but relatively small area of research suggests that poor or degrading data quality decreases the performance of AI models (see Zha et al., 2023 for a recent survey). Experimental studies show that dynamic changes in the distribution of training data can degrade algorithmic performance (related concepts in Machine Learning are concept drift, recursion, and model collapse). For example, in an exercise comparable to our empirical setting, Cai et al. (2021) and Jain and Shenoy (2023) study the computer vision problem of geolocalization. In essence, they test the prediction accuracy of a model that learns the geographic location depicted in photos, using both image data as well as geolocation data from sensors in the camera device. Resembling the example of a changing city skyline discussed above, the authors show that the prediction accuracy of a model trained on a batch of data from a particular time period is highest when images in the test dataset are from the same or previous ranges, but accuracy quickly declines when images in the test dataset are from subsequent periods. The natural shift in the distribution of images implies a reduction in model accuracy of about 4–9% over three years (see Figure 1 in Jain and Shenoy, 2023). Based on this, we can speculate that a reduction in novelty of about 13% to 28% beyond the natural rate, as we have estimated above, would translate into a decrease in output quality of 4.5% to 11.5%.
A potential solution is to impute missing or low-quality data with generated training data, i.e. synthetic data, to solve the issues of data quantity and quality (Lucini, 2021). Recent papers in law (e.g., Lee, 2024) and computer science (e.g. Gerstgrasser et al., 2024) discuss the potential for and address the concerns related to synthetic data and AI performance. However, a decrease in the variety of user-generated data, as in our empirical context, could still be detrimental to AI model performance.
A recent study points out how the injection of synthetic data into training data may not degrade AI outputs as long as it adds to, rather than replaces, the stock of data (see e.g., Gerstgrasser et al., 2024). However, assuming that the value of past data can decay over time, with rates depending on the use case or topic of interest (Valavi et al., 2022a), adding synthetic data effectively does replace the stock of training data over time. Thus, if we interpret AI-generated, synthetic, content as information similar to the input data, we can use the empirical estimates in this literature to speculate how a reduction in the variety in the flow of human-generated data translates into algorithmic performance. Experimental evidence suggests that adding AI-generated data to the training data implies a reduction in “accuracy” of 5% to 22% (see Figure 3 in Gerstgrasser et al., 2024, when Cross-Entropy Loss is interpreted as a measure of accuracy for the sake of a simple argument). Assuming that AI-generated content is identical to input data, our estimate of a 5% increase in similarity would reduce accuracy by 0.25% to 1.4%.
In conclusion, proposals that point out the potential to train models with synthetic data to circumvent human input and thus copyright-related challenges (e.g., Lee, 2024) need to be interpreted with caution. In situations where the value of the stock of data decays over time, issues related to AI model performance and eventual collapse can’t be addressed by synthetic data (Valavi et al., 2022a; Cai et al., 2021; Jain and Shenoy, 2023; Shumailov et al., 2024). While the temporal dynamics of data are considered exogenous in computer science (concept drift), the social sciences recognize it as endogenous (Quinn and Gutt, 2025), subject to the behavior of and incentives for the humans that produce the information. Hence, policy solutions need to be guided by the insight – i.e. strategic individual and organizational behavior govern the supply, and therefore quality of training data and eventually AI output.
5 Discussion
Creators across various industries have exhibited comparable responses to AI models trained on their work, reinforcing the broader applicability of our findings. Authors have expressed outrage upon discovering that AI models were trained on their copyrighted books without consent or compensation. For instance, Australian publisher Black Inc asked its authors to sign agreements allowing their works to be used for AI training, leading to significant backlash from the literary community. [^27] See https://www.theguardian.com/books/2025/mar/05/black-inc-melbourne-publisher-ai-agreements-writers-anger. Musicians have also raised concerns about AI-generated songs potentially diluting their revenue streams and reducing incentives for new compositions. [^28] See https://www.billboard.com/pro/labels-artists-sign-petition-opposing-ai-training-creative-works/. Over 1,000 artists released a silent album titled “Is This What We Want?” in protest against UK government proposals to allow AI companies to use copyright-protected works without permission. [^29] See https://www.theguardian.com/technology/2025/feb/25/kate-bush-damon-albarn-1000-artists-silent-ai-protest-album-copyright. A controversy on Instagram unfolded when artists learned their posts might be used to train Meta’s AI models, causing some users to migrate to platforms promising not to engage in AI data mining. In response, Meta has faced pressure to clarify or revise its policies regarding AI’s use of user content. [^30] See https://www.washingtonpost.com/technology/2024/06/06/instagram-meta-ai-training-cara/.
In what follows, we discuss key implications for managers and policy that arise from our results .
5.1 Managerial implications
Realizing the value of their content for genAI training purposes, platforms are increasingly partnering with AI firms to license their content archives to AI companies. [^31] OpenAI signed several agreements with newspapers such as the Guardian or Newscorp. See https://openai.com/index/openai-and-guardian-media-group-launch-content-partnership/ and https://www.nytimes.com/2024/05/22/business/media/openai-news-corp-content-deal.html. However, decision-makers should be mindful of data-sharing terms, as misaligned or exploitative policies can drive contributors – and their content – elsewhere. Our findings show that user responses to the exploitation of their content in AI training meaningfully affects both the quantity and quality of contributions. From a managerial perspective, this raises questions about how platform operators and other industry stakeholders can capitalize on AI’s potential without sacrificing the supply of human-generated content.
One approach is to provide multiple levels of data licensing and usage rights that incentivize continued engagement. For example, in October 2022 Unsplash introduced Unsplash+, a paid subscription service that explicitly forbids using images for machine learning, AI, and biometric tracking technology, and offers contributors monetary compensation for permitted uses. [^32] See https://unsplash.com/plus/license and https://unsplash.com/blog/contribute-to-unsplash/. Similarly, Shutterstock, another stock photography company, announced a “contributor fund” to compensate artists whose work is integrated into generative AI datasets. [^33] See https://support.submit.shutterstock.com/s/article/Shutterstock-ai-and-Computer-Vision-Contributor-FAQ?language=en_US. Alternatively, platforms can implement “opt-out” tags for AI crawlers or introduce specialized “AI-ready” tiers, which give creators enhanced levels of control. In a similar vein, AI companies are acknowledging creators’ concerns. After being confronted with artist complaints and a series of impending lawsuits, OpenAI developed tools to give content creators more control, such as allowing artists to opt out of having their art used in future model training. [^34] See https://www.wired.com/story/openai-olive-branch-artists-ai-algorithms/. As of this writing, it remains unclear whether these attempts will prove fruitful. Hence, we need additional research into how platforms can profit from genAI’s appetite for data without deterring contributor participation.
5.2 Market-level implications
Our analysis shows that adding user-generated content in AI training datasets can lead to behavior that negatively impacts the quantity, variety, and novelty of contributions, thereby reducing AI output quality. If policymakers and industry stakeholders do not adequately address creators’ concerns – particularly around fair compensation and rights management – both rights holders and AI-driven innovation may suffer in the long run. One potential path forward could be an automated dynamic compensation model that pays royalties based on the usage and commercial success of AI models trained with human-created works. On the one hand, a usage-based remuneration system would compensate contributors whenever their works are used in the training phase of an AI model (see Wang et al., 2024 for a technical proposal), much like music royalties that are paid out for each song played. Further, our results show that more intensely treated users display a more pronounced reaction. This highlights that a one-size-fits-all approach to licensing human-generated work for AI training datasets can lead to unintended heterogeneous effects on the flow of training data. Consequently, a licensing scheme that ensures a continued flow of data may need to compensate more active contributors to a larger degree. On the other hand, outcome-based bonuses could reward contributors if AI models built on their data achieve commercial success or lead to significant technological breakthroughs – akin to receiving royalties when a song is reused. Such a dual compensation mechanism would ensure that contributors are rewarded not just for the use of their data but also for the value it generates. Deploying such a system at scale would require advanced tracking technologies (e.g., digital fingerprinting) from training to output generation, as well as trusted intermediaries to manage potentially complex and large-scale transactions across various stakeholders in the AI ecosystem.
The existing data economy around the online advertising industry can serve as inspiration. Data auctions and data exchange platforms may provide efficient solutions. Data auctions facilitate efficient price discovery as AI developers and firms bid for the rights to use datasets, reflecting real-time market supply and demand. This competitive bidding process incentivizes creators to produce diverse, high-quality data likely to attract higher bids. The transparent nature of auctions provides clear benchmarks for data valuation, enhancing overall market transparency. Simultaneously, data exchange platforms could serve as centralized markets similar to stock exchanges but dedicated to trading data rights. Here, data creators can list their datasets, allowing buyers to purchase rights for specific uses. Much like rights collection societies in music, authorship, and the visual arts, these platforms would standardize transactions, reduce the complexity of individual agreements, and help ensure regulatory compliance. This may increase the liquidity of data as a commodity, making it accessible for AI developers when needed and expanding the market reach for data creators. By making pricing, supply, and demand visible, such platforms could enhance market efficiency, allowing buyers and sellers to make informed decisions based on up-to-date market information.
For policymakers, finding the right solution to the ongoing discussions surrounding AI training data and copyright will be crucial. Failure to do so could potentially lead to adverse outcomes, such as AI firms relocating their model training operations to copyright-friendly jurisdictions like Israel or Japan (Fiil-Flynn et al., 2022). Given the global competition in the regulation of digital markets, countries compete not only through tax incentives but also by offering attractive legal frameworks for AI research and development.